OpenStudy (anonymous):

Please Help!! Running out of time!!! Compare the follow sets of numbers to determine whether they are greater than, less than, or equal to. 1. the square root of 72 to 8 7/10 2. the square root of 97 to 9 2/11 3. the square root of 35 to 5 1/4 4. the square root of 50 to 7 4/5

OpenStudy (whpalmer4):

Some hints: \[\sqrt{72} = \sqrt{2*36}\]\[\sqrt{2}\approx 1.414\]